OpenStudy (anonymous):

Charlie was driving his boat at 30 mph for two hours against the current. He turned around and drove his boat the same speed but this time he was going with the current and it only took him and hour. How fast was he going in still water? Does this make sense?

OpenStudy (valpey):

Does this make sense? The part I am confused by is does the 30mph refer to boat speed minus current speed? If so then we have: Let c = current speed in mph Let b = boat speed in mph 30mph = b - c (mph) Distance = 30mph * 2 hours = 60 miles 60 miles = (b + c) * 1 hour 60 = b + c = 2*30 = 2*(b-c) b + c = 2b - 2c 3c = b 60 = 3c + c = 4c c = 15 mph b = 3c = 45 mph

OpenStudy (anonymous):

Thanks